OpenStudy (anonymous):

solve log (x-1) + Log 4=2

OpenStudy (anonymous):

You can rewrite sum of two logs as one log by multiplying the function

OpenStudy (anonymous):

Ok, umm, so you would multiply the (x-1) and the 4?

OpenStudy (anonymous):

(x-1)4=100 4x-4=100 x=(100+4)/4 x= 26

OpenStudy (anonymous):

wow! How did u get 100?

OpenStudy (anonymous):

\[\text{Log}[a]+\text{Log}[b]=\text{Log}[a*b]\]

OpenStudy (anonymous):

U do not go (x-1)4= 4x-4? Where did the 100 come in?

OpenStudy (anonymous):

Let me walk through you, based on my post above \[\text{Log}[(x-1)(4)]=2\] \[\text{Log}[(4x-4)]=2\] Now change that to exponential form, \[\text{Log}_ba=c\space \text{ }\space \text{--}> b^c=a\] 10^2=4x-4 100=4x-4

OpenStudy (anonymous):

i will look and see if i got it

OpenStudy (anonymous):

ok, that part makes sense. so the next step is to...?

OpenStudy (anonymous):

oh, i see it.

OpenStudy (anonymous):

Thank you! God bless